# Convergence of Schrödinger operators on domains with scaled resonant potentials

### v1

@@ 3.3. Proof of Theorem A
$\qquad\qquad+\big{(}u,({\mathsf{I}}-{\mathsf{J}}_{\varepsilon}){\mathsf{R}}_{\varepsilon}^{*}v\bigr{)}-\bigl{(}({\mathsf{I}}-{\mathsf{J}}_{\varepsilon}){\mathsf{R}}u,v\bigr{)}$
@@
using first representation theorem for the sesquilinear forms $\mathfrak{a}$ and $\mathfrak{a}_{\varepsilon}$ associated with ${\mathsf{A}}$ and ${\mathsf{A}}_{\varepsilon}$, respectively, and using also the self-adjointness of ${\mathsf{I}}-{\mathsf{J}}_{\varepsilon}$ for the last equality. Moreover, we have $\lVert{\mathsf{R}}u\rVert\leq\lVert u\rVert$ and $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert\leq\lVert v\rVert$ as ${\mathsf{R}}$ and ${\mathsf{R}}_{\varepsilon}^{*}$ are the resolvents at the points ±$\pm{\mathsf{i}}$ and ${\mathsf{A}}$ and ${\mathsf{A}}_{\varepsilon}$ are self-adjoint, respectively, hence we conclude (using Cauchy-Schwarz)
$\big{\lvert}\big{\langle}({\mathsf{R}}_{\varepsilon}-{\mathsf{R}})u,v\big{\rangle}\big{\rvert}$ $\leq\big{\lvert}\mathfrak{a}[{\mathsf{R}}u,{\mathsf{J}}_{\varepsilon}{\mathsf{R}}_{\varepsilon}^{*}v]-\mathfrak{a}_{\varepsilon}[{\mathsf{J}}_{\varepsilon}{\mathsf{R}}u,{\mathsf{R}}^{*}_{\varepsilon}v]\big{\rvert}+\big{\lVert}({\mathsf{J}}_{\varepsilon}-{\mathsf{I}})u\big{\rVert}\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert+\big{\lVert}({\mathsf{J}}_{\varepsilon}-{\mathsf{I}}){\mathsf{R}}u\big{\rVert}\lVert v\rVert$
@@
Now all norms on $\Omega_{a\varepsilon}$ converge to 00 by Lebesgue’s convergence theorem for $u\in C^{\infty}_{\mathrm{c}}(\Omega)$. By density of $C^{\infty}_{\mathrm{c}}(\Omega)$ in $L_{2}(\Omega)$, we conclude that ${\mathsf{A}}_{\varepsilon}$ converges to ${\mathsf{A}}$ in strong resolvent sense.
4. Convergence for non-negative potentials
@@ 4.1. Auxiliary boundary mappings, the identification operator and some related estimates
Recall that we fix $a>0$ such that $\operatorname{supp}V\subset[0,a]$. Let $\psi_{0}$ be a non-resonant solution of (1.1) normalised so that $\psi_{0}^{\prime}(x)=1$ for all $x>a$; see Remark 1.2 (b).
@@
[Proof]The proof is exactly the same as the proof of Lemma 3.3 — except that by differentiating the identity $-\psi_{0}^{\prime\prime}+V\psi_{0}=0$ we obtain $-\psi_{0}^{\prime\prime\prime}+V^{\prime}\psi_{0}+V\psi_{0}^{\prime}=0$. In particular, the first and fourth term in (3.4) do not cancel, as now $\psi_{0}^{\prime\prime\prime}-V\psi_{0}^{\prime}=V^{\prime}\psi_{0}$ remains. In particular, we have
@@
from which the desired formula follows. Here the boundary term at $t=0$ vanishes due to $\widetilde{v}(s,0)=0$ while the boundary terms at $t=a\varepsilon$ vanishes due to $\psi_{0}^{\prime\prime}(a)=0$.
@@
$\big{\lvert}\mathfrak{a}_{0}[u,{\mathsf{K}}_{\varepsilon}v]-\mathfrak{a}_{\varepsilon}[{\mathsf{K}}_{\varepsilon}u,v]\big{\rvert}\leq\hat{c}_{0}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}$
@@
[Proof]We estimate the first term in Lemma 4.4 as in the proof of Lemma 3.4 writing $\hat{c}^{\prime}$ for the constant as in (3.5) with $\psi_{0}^{\prime}$ replaced by $\psi_{0}^{\prime\prime}$); this gives
@@
$\leq\hat{c}^{\prime}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{L_{2}(\Omega_{a\varepsilon})}+\frac{\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}}{\varepsilon^{2}}\int_{0}^{a\varepsilon}\lVert\Gamma_{t}u\rVert_{L_{2}(\Sigma)}\lVert\widetilde{\Gamma}_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq\hat{c}^{\prime}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{L_{2}(\Omega_{a\varepsilon})}+\frac{\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{\rho\varepsilon^{2}}\int_{0}^{a\varepsilon}\!\!t\operatorname{d\!}t\lVert\nabla u\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}\lVert\nabla v\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}$
$=\hat{c}^{\prime}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{L_{2}(\Omega_{a\varepsilon})}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\lVert\nabla u\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}\lVert\nabla v\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}$
$\leq\Bigl{(}\hat{c}^{\prime}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\Bigr{)}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}$
$\leq\hat{c}_{0}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}\qquad\text{where}\qquad\hat{c}_{0}\vcentcolon=\Bigl{(}\hat{c}^{\prime}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\Bigr{)}$ (4.2)
@@ 4.2. Proof of Theorem B
Before providing the proof of Theorem B, we need one more estimate: here, it is essential that the potential is non-negative:
@@
Assume that $V\geq 0$ then we have
$\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega)}\leq\sqrt{2}\lVert v\rVert$
for $v\in L_{2}(\Omega)$.
[Proof]The estimate $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{L_{2}(\Omega)}\leq\lVert v\rVert$ is clear by spectral calculus. Moreover, we have for $\hat{v}\in\operatorname{dom}{\mathsf{A}}_{\varepsilon}$
$\lVert\nabla\hat{v}\rVert_{L_{2}(\Omega;{\mathbb{C}}^{d})}^{2}$ $\leq\lVert\nabla\hat{v}\rVert_{L_{2}(\Omega;{\mathbb{C}}^{d})}^{2}+\Big{\langle}V_{\varepsilon}\hat{v},\hat{v}\Big{\rangle}=\mathfrak{a}_{\varepsilon}[\hat{v}]$
$=\langle{\mathsf{A}}_{\varepsilon}\hat{v},\hat{v}\rangle\leq\lVert{\mathsf{A}}_{\varepsilon}\hat{v}\rVert^{2}+\lVert\hat{v}\rVert^{2}=\lVert({\mathsf{A}}_{\varepsilon}+{\mathsf{i}})\hat{v}\rVert^{2}$
as $V\geq 0$ and using the spectral calculus stating that ${\mathsf{A}}_{\varepsilon}\leq{\mathsf{A}}_{\varepsilon}^{2}+{\mathsf{I}}$ in the form sense for the last inequality. The desired estimate follows by setting $\hat{v}={\mathsf{R}}_{\varepsilon}^{*}v$.
@@
$\big{\lvert}\langle({\mathsf{R}}_{\varepsilon}-{\mathsf{R}}_{0})u,v\rangle\big{\rvert}$
@@
$\leq\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl{(}\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\lVert v\rVert\bigr{)}$
$\leq\Bigl{(}\sqrt{2}\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{2}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl{(}\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\bigr{)}\Bigr{)}\lVert v\rVert$
using Lemma 4.3 and Lemma 4.5 in the second estimate, and $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega)}\leq\sqrt{2}\lVert v\rVert$ resp. Lemma 4.6 in the last one. We conclude that
@@
$\leq\sqrt{2}\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{2}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl{(}\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\bigr{)}.$
As before, all norms on $\Omega_{a\varepsilon}$ converge to 00 by Lebesgue’s convergence theorem for $u\in C^{\infty}_{\mathrm{c}}(\Omega)$. as $\varepsilon\to 0$. By density of $C^{\infty}_{\mathrm{c}}(\Omega)$ in the respective spaces, we conclude that ${\mathsf{A}}_{\varepsilon}$ converges to ${\mathsf{A}}_{0}$ in the strong resolvent sense.

### v2

@@ 3.3. Proof of Theorem A
using first representation theorem for the sesquilinear forms $\mathfrak{a}$ and $\mathfrak{a}_{\varepsilon}$ associated with $\mathsf{A}$} and ${\mathsf{A}}_{\varepsilon}$, respectively, and using also the self-adjointness of ${\mathsf{I}}-{\mathsf{J}}_{\varepsilon}$ for the last equality. Moreover, we have $\lVert{\mathsf{R}}u\rVert\leq\lVert u\rVert$ and $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert\leq\lVert v\rVert$ as $\mathsf{R}$} and ${\mathsf{R}}_{\varepsilon}^{*}$ are the resolvents at the points $\pm{\mathsf{i}}$ and $\mathsf{A}$} and ${\mathsf{A}}_{\varepsilon}$ are self-adjoint, respectively, hence we conclude (using Cauchy-Schwarz)
[…]
@@
Now all norms on $\Omega_{a\varepsilon}$ converge to 0 by Lebesgue’s convergence theorem for $u\in C^{\infty}_{\mathrm{c}}(\Omega)$. By density of $C^{\infty}_{\mathrm{c}}(\Omega)$ in $L_{2}(\Omega)$, we conclude that ${\mathsf{A}}_{\varepsilon}$ converges to $\mathsf{A}$} in strong resolvent sense.
4. Convergence for non-resonant potentials
@@ 4.1. Auxiliary boundary mappings, the identification operator and some related estimates
Recall that we fix $a>0$ such that $\operatorname{supp}V\subset[0,a]$. Let $\psi_{0}$ be a non-resonant solution of (1.1) normalised so that $\psi_{0}^{\prime}(t)=1$ for all $t>a$; see Remark 1.2 (b).
@@
[Proof]The proof is exactly the same as the proof of Lemma 3.3 — except that by differentiating the identity $\psi_{0}^{\prime\prime}+V\psi_{0}=0$ we obtain $\psi_{0}^{\prime\prime\prime}+V^{\prime}\psi_{0}+V\psi_{0}^{\prime}=0$. In particular, the first and fourth term in (3.4) do not cancel, as now $\psi_{0}^{\prime\prime\prime}-V\psi_{0}^{\prime}=V^{\prime}\psi_{0}$ remains. In particular, we have
@@
from which the desired formula follows. Here the boundary term at $t=0$ vanishes due to $\widetilde{v}(s,0)=0$ while the boundary term at $t=a\varepsilon$ vanishes due to $\psi_{0}^{\prime\prime}(a)=0$.
@@
$\big\lvert\mathfrak{a}_{0}[u,{\mathsf{K}}_{\varepsilon}v]-\mathfrak{a}_{\varepsilon}[{\mathsf{K}}_{\varepsilon}u,v]\big\rvert\leq\hat{c}_{0}\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}$
@@
[Proof]We estimate the first term in Lemma 4.4 as follows
$\bigg|\frac{1}{\varepsilon}$ $\int_{0}^{a\varepsilon}\psi_{0}^{\prime\prime}\Bigl(\frac{t}{\varepsilon}\Bigr)\big\langle\Upsilon_{t}u,\Gamma_{t}v\big\rangle_{L_{2}(\Sigma)}\operatorname{d\!}t\bigg|$
$\!\leq\frac{\|\psi_{0}^{\prime\prime}\|_{\infty}}{\varepsilon}\left(\int_{0}^{a\varepsilon}\|\Upsilon_{t}u\|^{2}_{L_{2}(\Sigma)}\operatorname{d\!}t\right)^{1/2}\left(\int_{0}^{a\varepsilon}\|\Gamma_{t}v\|^{2}_{L_{2}(\Sigma)}\operatorname{d\!}t\right)^{1/2}$
$\!\leq\frac{a\|\psi_{0}^{\prime\prime}\|_{\infty}\|v\|_{H^{1}(\Omega_{a\varepsilon})}}{\sqrt{2\rho}}\left(\int_{0}^{a\varepsilon}\int_{\Sigma}\left(8\varphi^{2}(s,t)|\partial_{t}\widetilde{u}(s,t)|^{2}+2(\partial_{t}\varphi(s,t))^{2}|\widetilde{u}(s,t)|^{2}\right)\operatorname{d\!}\sigma(s)\operatorname{d\!}t\right)^{1/2}$
$\!\leq\frac{2a\|\psi_{0}^{\prime\prime}\|_{\infty}\|v\|_{H^{1}(\Omega_{a\varepsilon})}}{\rho}\max\{\|\varphi\|_{\infty},\|\partial_{t}\varphi\|_{\infty}\}\!\left(\int_{0}^{a\varepsilon}\!\int_{\Sigma}\left[|\partial_{t}\widetilde{u}(s,t)|^{2}+|\widetilde{u}(s,t)|^{2}\right]\varphi(s,t)\operatorname{d\!}\sigma(s)\operatorname{d\!}t\right)^{1/2}$
$\!\leq\hat{c}^{\prime}\|u\|_{H^{1}(\Omega_{a\varepsilon})}\|v\|_{H^{1}(\Omega_{a\varepsilon})},\qquad\quad\text{where}\qquad\hat{c}^{\prime}:=\frac{2a\|\psi_{0}^{\prime\prime}\|_{\infty}}{\rho}\max\big\{\|\varphi\|_{\infty},\|\partial_{t}\varphi\|_{\infty}\big\},$
where we used Cauchy-Schwarz inequality in the first estimate and Lemma 4.1 in the second. Employing the above estimate we get
@@
$\leq\hat{c}^{\prime}\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}+\frac{\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}}{\varepsilon^{2}}\int_{0}^{a\varepsilon}\lVert\Gamma_{t}u\rVert_{L_{2}(\Sigma)}\lVert\widetilde{\Gamma}_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq\hat{c}^{\prime}\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}+\frac{\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{\rho\varepsilon^{2}}\int_{0}^{a\varepsilon}\!\!t\operatorname{d\!}t\lVert\nabla u\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}\lVert\nabla v\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}$
$=\hat{c}^{\prime}\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\lVert\nabla u\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}\lVert\nabla v\rVert_{L_{2}(\Omega_{a\varepsilon};{\mathbb{C}}^{d})}$
$\leq\Bigl(\hat{c}^{\prime}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\Bigr)\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}$
$\leq\hat{c}_{0}\lVert u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert v\rVert_{H^{1}(\Omega_{a\varepsilon})}\qquad\text{where}\qquad\hat{c}_{0}\vcentcolon=\Bigl(\hat{c}^{\prime}+\frac{a^{2}\lVert V^{\prime}\rVert_{\infty}\lVert\psi_{0}\rVert_{L_{\infty}((0,a))}\lVert\varphi\rVert_{\infty}}{2\rho}\Bigr)$ (4.2)
@@ 4.2. Proof of Theorem B
Before providing the proof of Theorem B, we need one more estimate: here, it is essential that the potential satisfies an assumption on the negative part:
@@
Let the constant $C_{\Omega}>0$ be defined as in (1.7). Under the assumption
$C_{V}:=\max_{t\in{\mathbb{R}}_{+}}(-V(t)t^{2})<C_{\Omega}$ (4.3)
we have
$\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega)}\leq\sqrt{c_{\Omega,V}+1}\lVert v\rVert$
for all $v\in L_{2}(\Omega)$ and the constant $c_{\Omega,V}:=\frac{1}{2}\big(1-\frac{C_{V}}{C_{\Omega}}\big)^{-1}\geq\frac{1}{2}$.
[Proof]The estimate $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{L_{2}(\Omega)}\leq\lVert v\rVert$ is clear by spectral calculus. Under the assumption we stated on the potential $V$, we have the inequality
$V_{\varepsilon}(x)\geq-\frac{C_{V}}{({\rm dist}\,(x,\Sigma))^{2}},$
for all $\varepsilon>0$ and all $x\in\Omega$. We have by the definition of the constant $C_{\Omega}$ that for all $\hat{v}\in\operatorname{dom}{\mathsf{A}}_{\varepsilon}$
$\lVert\nabla\hat{v}\rVert_{L_{2}(\Omega;{\mathbb{C}}^{d})}^{2}$ $\leq 2c_{\Omega,V}\left(\lVert\nabla\hat{v}\rVert_{L_{2}(\Omega;{\mathbb{C}}^{d})}^{2}+\Big\langle V_{\varepsilon}\hat{v},\hat{v}\Big\rangle\right)=2c_{\Omega,V}\mathfrak{a}_{\varepsilon}[\hat{v}$]
$=2c_{\Omega,V}\langle{\mathsf{A}}_{\varepsilon}\hat{v},\hat{v}\rangle\leq c_{\Omega,V}\left(\lVert{\mathsf{A}}_{\varepsilon}\hat{v}\rVert^{2}+\lVert\hat{v}\rVert^{2}\right)=c_{\Omega,V}\lVert({\mathsf{A}}_{\varepsilon}+{\mathsf{i}})\hat{v}\rVert^{2}$
and using the Cauchy-Schwarz inequality combined with standard inequality $2ab\leq a^{2}+b^{2}$ for $a,b\geq 0$ in the penultimate step. The desired estimate follows by setting $\hat{v}={\mathsf{R}}_{\varepsilon}^{*}v$.
@@
$\big\lvert\langle({\mathsf{R}}_{\varepsilon}-{\mathsf{R}}_{0})u,v\rangle\big\rvert\hskip-43.05542pt$
@@
$\leq\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{1}(\Omega_{a\varepsilon})}\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl(\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\lVert v\rVert\bigr)$
$\leq\Bigl(\sqrt{c_{\Omega,V}+1}\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{1}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl(\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\bigr)\Bigr)\lVert v\rVert$
using Lemma 4.3 and Lemma 4.5 in the second estimate, and $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega)}\leq\sqrt{c_{\Omega,V}+1}\lVert v\rVert$ resp. Lemma 4.6 in the last one. We conclude that
@@
$\leq\sqrt{c_{\Omega,V}+1}\hat{c}_{0}\lVert{\mathsf{R}}_{0}u\rVert_{H^{1}(\Omega_{a\varepsilon})}+\lVert\psi_{0}^{\prime}-1\rVert_{\infty}\bigl(\lVert u\rVert_{L_{2}(\Omega_{a\varepsilon})}+\lVert{\mathsf{R}}_{0}u\rVert_{L_{2}(\Omega_{a\varepsilon})}\bigr).$
As before, all norms on $\Omega_{a\varepsilon}$ converge to 0 by Lebesgue’s convergence theorem for $u\in C^{\infty}_{\mathrm{c}}(\Omega)$. as $\varepsilon\to 0$. By density of $C^{\infty}_{\mathrm{c}}(\Omega)$ in the respective spaces, we conclude that ${\mathsf{A}}_{\varepsilon}$ converges to ${\mathsf{A}}_{0}$ in the strong resolvent sense.